# Yuri Balashov

### v1

@@ 3.1 Model Architecture and Size
Typical state-of-the-art NMT models have 100M to 10B parameters and 6–12 transformer layers on the encoder and decoder sides. In contrast, cutting-edge LLMs are rumored to have trillions of parameters (two to three orders of magnitude larger). NMT models are usually encoder-decoder transformers: an encoder processes the source sentence, and a decoder generates the target sentence, attending to the encoder’s representation. Most LLMs, on the other hand, are decoder-only transformers, especially those designed for text generation. A decoder-only model produces output based on a single stream of input, which can include instructions or source text. Despite this difference, an LLM can effectively act like an encoder-decoder when the prompt includes the source text: the instruction implicitly tells it to “encode” the source meaning into its latent state before producing the output; for example: Translate the following sentence to French: “I hate squirrels.”
@@ 3.2 Training Data
This is perhaps the starkest difference. An NMT model is trained on a relatively narrow and specific dataset: a large collection of parallel corpora (source-target sentence pairs). For example, a high-quality English-German NMT model might be trained on tens of millions of aligned sentence pairs from sources like EU Parliament proceedings, translated news, subtitles, and so forth. The NMT model thus directly learns to map between two languages. In contrast, an LLM is typically trained on a broad swath of text scraped from the Internet, predominantly in one language (most often English) with a mix of others, and not specifically on parallel data. For instance, GPT-style models are trained on web pages, books, and articles comprising many billions of words, of which only a small fraction may be parallel or even multilingual content (Blevins and Zettlemoyer, 2022; Chowdhery et al., 2022). As a result, NMT systems have in-domain strength for translation but limited “knowledge” outside language mappings, whereas LLMs boast vast “knowledge” and multilingual exposure but only implicit translation mappings.
@@ 3.3 Training Objective
The dedicated NMT’s training objective is explicitly to maximize translation likelihood, essentially directly training to translate. The loss function compares the model’s output with reference translations for each source sentence. The LLM’s objective, by contrast, is next-word prediction (language modeling) across its training texts. No direct signal tells the LLM “this is a translation of that.” Instead, at best, some training documents might include translations or multilingual sentence collections, and the LLM learns to predict those sequences. In short, translation is the end goal for NMT but a mere side effect for LLMs. This difference is crucial: LLMs do not optimize for fidelity or adequacy explicitly in translation, which is why they may sometimes prioritize fluent output over precise correspondence (a known issue with LLM translations, according to human evaluations).
@@ 3.4 Inference and Use
Using an NMT model typically means feeding a source sentence into the model’s encoder and letting the decoder produce a translation. It is a straightforward, single-shot process. As already noted, using an LLM for translation requires crafting a suitable prompt. For example, to translate from English to Spanish, one might prompt the LLM with: Translate the following English sentence to Spanish: <source sentence>. The LLM then generates the Spanish text as a continuation. The quality of the LLM’s translation can be sensitive to prompt wording (Jiao et al., 2023b; Vilar et al., 2023; Peng et al., 2023). The presence of examples (few-shot prompting; more on this below) or specifying the desired style can change results. By contrast, an NMT model’s output is largely deterministic given the source. This means that LLMs offer more flexibility (one can prompt for a literal translation or a creative one, or ask an LLM to translate and explain), but also more variability. LLMs might ignore the instruction if not phrased clearly or might produce extra text (like apologies or analysis) if the prompt isn’t tight (Zhang et al., 2023a; Peng et al., 2023; Lyu et al., 2023; Balashov et al., 2025).
@@ 3.5 Quality and Evaluation
How do these approaches compare in translation quality? Recent studies have benchmarked LLMs against dedicated MT systems. For example, Zhang et al. (2023b) evaluated 15 open-source language models on translation tasks via prompting and compared to fine-tuned NMT models. They found that a bit of fine-tuning or domain adaptation (e.g. via QLoRA, Dettmers et al. 2023) on a small translation dataset boosts the translation performance of the model very significantly. A more extensive study showed that GPT-4 (which was a leading model at that time) outperformed Meta’s NLLB (No Language Left Behind, a specialized multilingual NMT model, Team et al. 2022) in about 40% of evaluated translation directions, but still underperformed a strong commercial engine (Google Translate) on many low-resource language pairs (Zhu et al., 2024). The consensus emerging from the work reported in 2023–2024 was that LLMs are remarkably good general translators, especially for high-resource languages and with careful prompting, but a dedicated NMT that is tailored to a specific domain or style can still have an edge in its niche. For example, for technical texts with specialized terminology, a fine-tuned NMT might preserve terminology more consistently than an LLM, unless the LLM is explicitly guided or instructed (Hendy et al., 2023). But the field is developing very rapidly. [^5] For a recent industry report, see Intento, Inc. 2025. For a recent linguistically-focused comparison of NMT, LLM, and human translation outputs, see Sizov et al. 2024.
@@ 3.6 Multilingualism
While some recent NMT models are multilingual by design (i.e. trained on parallel sentences in multiple languages and capable of translating between a chosen pair; e.g. NLLB), most are bilingual (one language pair). LLMs, in contrast, tend to be trained on multiple languages as a necessity: even if 90–95% of the data is English, the remaining 5–10% might cover dozens of languages (Blevins and Zettlemoyer, 2022; Meta, 2025). As a result, LLMs often have broader multilingual scope than a given NMT model. A single LLM like GPT-4o or Llama can translate between many languages (English to French, French to German, Japanese to Russian, etc.), whereas an NMT model is usually limited to the directions it was trained on (unless it’s a massively multilingual model like NLLB). This breadth is advantageous, but it comes with uncertainty: the LLM may not be equally strong in all languages. Some languages with very little representation in pretraining data see much lower translation quality. [^6] E.g., Zhu et al. (2024) note that GPT-4 still struggled on several African and Southeast Asian languages. NMT models can also struggle with low-resource languages, of course, but one can explicitly gather parallel data for those languages to train or fine-tune the model, whereas an LLM’s fixed pretraining mix might under-represent them (unless augmented with further training).
@@ 3.7 Behavior and Errors
NMT and LLM systems also tend to exhibit different failure modes. NMT models, when they err, often produce mistranslations that are locally implausible (e.g. garbled or nonsensical phrases) or they might drop segments of the input (omissions) if attention-based alignment fails. LLMs, by contrast, rarely produce outright garble; thanks to their strong language modeling priors their errors are usually fluent but can be misleading. A common LLM translation error is hallucination—inserting content absent from the source, or over-smoothing—translating idiomatic expressions literally, thus losing nuance. [^7] But He et al. (2024) report that prompting an LLM to mimic a human translator’s step-by-step process can reduce such deficiencies. LLMs might also be inconsistent in style or level of formality unless instructed, whereas an NMT system can be constrained by training data style or use explicit formality tags. These differences remind us that an LLM is trying to produce a plausible continuation given the prompt and its training, whereas an NMT is laser-focused on reproducing the source content in target form.
@@ 4 Recent Uses of LLMs in Translation Tasks
The application of LLMs to translation-related tasks began almost as soon as LLMs themselves burst onto the scene (Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023). Researchers and practitioners have tested cutting-edge LLMs on various multilingual tasks, often with impressive results. In this section, I review recent uses of LLMs in translation and other cross-lingual contexts. Key themes include zero-shot vs. few-shot prompting, performance on low-resource languages, pivot translation strategies, chain-of-thought prompting for translation, and exploiting the inherently multilingual nature of latent representations in LLMs.
@@ 4.1 Zero- and Few-Shot Translation with LLMs
One of the headline capabilities of LLMs is zero-shot learning—performing a task without direct task-specific training (Brown et al., 2020). Translation has become a primary example. Prompting an LLM with a simple instruction “Translate the following sentence to Mandarin:________” can yield a credible translation even if the model was never explicitly trained on parallel data for that language. Early tests with GPT-3 showed surprisingly competent zero-shot translation for several languages, sparking immediate interest in using such models as universal translators (Savenkov, 2023).
Subsequent research has systematically evaluated zero-shot vs. few-shot prompting for translation (e.g. Zhang et al. 2023b). Few-shot prompting means providing a few example translations in the prompt before asking the model to translate a new sentence. For instance, one might input:
@@
and let the model continue. This technique often boosts translation accuracy, especially for models not specifically instruction-tuned. Garcia et al. (2023) dubbed it the “unreasonable effectiveness of few-shot learning,” finding that even 1–2 examples can significantly improve the translation output output for certain LLMs. The improvement is typically more pronounced for smaller or less instruction-savvy models. On the other hand, for GPT-4 and later models, the gap between zero-shot and few-shot regimes is smaller, since the former is already strong. Later work noted that few-shot translation prompting can, in effect, make up for the lack of translation-specific instruction tuning, or of any instruction tuning.
In fact, curious phenomena have been observed. For example, Zhu et al. (2024) found that, when given in-context exemplars, LLMs sometimes ignore “unreasonable instructions” asking the model to translate into a different language or do something else altogether, e.g. summarize a sentence instead of translating it. The authors also found that cross-lingual exemplars (in different language pairs than the target pair) can sometimes provide better guidance for low-resource translation than same-pair exemplars. Thus, inserting several RU-EN sentence pairs in the prompt, following by un unrelated Chinese sentence produces a better English translation of it than ZH-EN examples. A hypothetical explanation may be that showing the model examples of, say, RU-EN translation may activate its “translation mode” generally, even if we actually want it to translate from Chinese to English, a direction for which it saw no direct examples. This aligns with the idea that the LLM has a general concept of “translating” that is somewhat language-agnostic. The general lesson appears to be that instruction semantics can be overridden by example context.
@@ 4.2 Low-Resource Languages
For extremely low-resource languages, with virtually no representation in the LLM’s pre-training data, neither pivoting nor prompt optimization may suffice. In such cases, researchers have explored fine-tuning LLMs on supplemental parallel data in a low-resource language pair. The hope is the LLM’s general knowledge transfers, and only minimal data is needed to teach it a new language mapping. Early results show some promise, but the cost of fine-tuning an LLM can be prohibitive. Fortunately, techniques such as LoRA (Low-Rank Adaptation, Hu et al. 2021) can be used to inject translation capabilities for new languages into an existing LLM at relatively low compute cost. Thus Zhang et al. (2023b) used an even more efficient method (QLoRA, Dettmers et al. 2023) to effectively turn a general LLM into a specialized translator with just a fraction of parameters updated, something that can be done on a single consumer-grade GPU.
@@ 4.3 Pivoting and Intermediate Languages
Pivoting (translating through an intermediate language) has been a longstanding strategy in MT for dealing with language pairs lacking direct parallel data. With LLMs, explicit pivoting can be attempted by prompting the model to go through two steps; for example: “Translate this sentence from German to English first; then translate the English output to Hindi.” This prompt can sometimes yield a better Hindi translation than prompting the model to translate directly from German to Hindi (cf. Jiao et al. 2023b).
@@ 4.4 Chain-of-Thought Prompting for Translation
Chain-of-thought (CoT) prompting is a technique where the model is asked to produce intermediate reasoning steps before the final answer (Wei et al., 2022b). Originally developed for math and logic problems, it has been applied to help LLMs improve their translation output by following, in some cases, a human translation workflow model. For example, Briakou et al. (2024) engaged Gemini 1.5 Pro in a multi-turn interaction involving pre-translation research, drafting, refining, and proofreading, which resulted in notable quality improvements.
CoT prompting could be particularly useful when dealing with longer documents. He et al. (2024) adopted this approach and asked LLMs to generate keywords, topics, and potential translations of tricky terms before attempting the full translation and then used those to guide the process. Since state-of-the-art LLMs have very large context windows (1M+ tokens in some cases), one can feed an entire document and prompt the model to work section by section. Researchers are actively exploring this: for instance, how to prompt an LLM to decide when to translate literally and when to adapt, essentially giving the model a strategy akin to a human translator’s decision-making (Jiao et al., 2023a; Lu et al., 2024a). This ventures into the territory of interactive translation with LLMs: rather than using a single prompt, one can craft a sequence of prompts for a genuinely collaborative scenario. There have been experiments using LLMs as translator assistants, where the human can query the model for alternate phrasings, explanations of source text nuances, error annotation, and more (Jiao et al., 2023a; Ki and Carpuat, 2024; Treviso et al., 2024).
Another interesting recent study develops “compositional translation,” a new method that uses LLMs to improve translation quality for low-resource languages (Zebaze et al., 2025). The approach works by breaking down complex sentences into simpler, semantically coherent phrases, translating each phrase using contextually similar examples, and then merging these translations into a coherent final result. Experiments show that this method can outperform traditional few-shot translation techniques in LLMs, making it a promising solution for low-resource challenges.
@@ 4.5 Multilingual Latent Representations and Cross-Lingual Tasks
Beyond direct translation, LLMs have shown capability in a variety of cross-lingual tasks, such as cross-lingual question answering (answering in English from a French context, etc.), multilingual summarization (summarizing a document in another language), code-switched dialogue understanding, and so on (Chua et al., 2023; Gao et al., 2025). Presumably, all these tasks leverage the model’s multilingual latent space. But what, exactly, is the shape of this space? Where in the model should we look for it? And what methods could be used to study it?

### v3

@@ 3.1 Model Architecture and Size
Typical state-of-the-art NMT models have 100M to 10B parameters and 6–12 transformer layers on the encoder and decoder sides. In contrast, cutting-edge LLMs are rumored to have trillions of parameters (two to three orders of magnitude larger). NMT models are usually encoder-decoder transformers: an encoder processes the source sentence, and a decoder generates the target sentence, attending to the encoder’s representation (Figure 2). Most LLMs, on the other hand, are decoder-only transformers, especially those designed for text generation. A decoder-only model produces output based on a single stream of input, which can include instructions or source text (Figure 3). Despite this difference, an LLM can effectively act like an encoder-decoder when the prompt includes the source text: the instruction implicitly tells it to “encode” the source meaning into its latent state before producing the output; for example: Translate the following sentence to Spanish: “My name is John.”
Figure 2: NMT architecture
Figure 3: LLM architecture
@@ 3.2 Training Data
This is perhaps the starkest difference. An NMT model is trained on a relatively narrow and specific dataset: a large collection of parallel corpora (source-target sentence pairs). For example, a high-quality English-German NMT model might be trained on tens of millions of aligned sentence pairs from sources like EU Parliament proceedings, translated news, subtitles, and so forth. The NMT model thus directly learns to map between two languages. In contrast, an LLM is typically trained on a broad swath of text scraped from the Internet, predominantly in one language (most often English) with a mix of others, and not specifically on parallel data. For instance, GPT-style models are trained on web pages, books, and articles comprising many billions of words, of which only a small fraction may be parallel or even multilingual content (blevins_language_2022; chowdhery_palm_2022). As a result, NMT systems have in-domain strength for translation but limited “knowledge” outside language mappings, whereas LLMs boast vast “knowledge” and multilingual exposure but only implicit translation mappings.
@@ 3.3 Training Objective
The dedicated NMT’s training objective is explicitly to maximize translation likelihood, essentially directly training to translate. The loss function compares the model’s output with reference translations for each source sentence. The LLM’s objective, by contrast, is next-word prediction (language modeling) across its training texts. No direct signal tells the LLM “this is a translation of that.” Instead, at best, some training documents might include translations or multilingual sentence collections, and the LLM learns to predict those sequences. In short, translation is the end goal for NMT but a mere side effect for LLMs.
@@ 3.4 Inference and Use
Using an NMT model typically means feeding a source sentence into the model’s encoder and letting the decoder produce a translation. It is a straightforward, single-shot process. As already noted, using an LLM for translation requires crafting a suitable prompt. For example, to translate from English to Spanish, one might prompt the LLM with: Translate the following English sentence to Spanish: <source sentence>. The LLM then generates the Spanish text as a continuation. The quality of the LLM’s translation can be sensitive to prompt wording (jiao_is_2023; vilar_prompting_2023; peng_towards_2023). The presence of examples (few-shot prompting; more on this below) or specifying the desired style can change results. By contrast, an NMT model’s output is largely deterministic given the source. This means that LLMs offer more flexibility (one can prompt for a literal translation or a creative one, or ask an LLM to translate and explain), but also more variability. LLMs might ignore the instruction if not phrased clearly or might produce extra text (like apologies or analysis) if the prompt isn’t tight (zhang_prompting_2023; peng_towards_2023; lyu_new_2023; balashov-etal-2025-translation).
@@ 3.5 Quality and Evaluation
How do these approaches compare in translation quality? Earlier studies benchmarked LLMs against dedicated MT systems. For example, zhang_machine_2023 evaluated 15 open-source language models on translation tasks via prompting and compared to fine-tuned NMT models. They found that a bit of fine-tuning or domain adaptation (e.g. via QLoRA, dettmers_qlora_2023) on a small translation dataset boosts the translation performance of the model very significantly. A more extensive study showed that GPT-4 (which was a leading model at that time) outperformed Meta’s NLLB (No Language Left Behind, a specialized multilingual NMT model, team_no_2022) in about 40% of evaluated translation directions, but still underperformed a strong commercial engine (Google Translate) on many low-resource language pairs (zhu_multilingual_2024). The consensus emerging from the work reported in 2023–2024 was that LLMs are remarkably good general translators, especially for high-resource languages and with careful prompting, but a dedicated NMT that is tailored to a specific domain or style can still have an edge in its niche. For example, for technical texts with specialized terminology, a fine-tuned NMT might preserve terminology more consistently than an LLM, unless the LLM is explicitly guided or instructed (hendy_how_2023). But the field is developing very rapidly. [^4] For a recent industry report, see intento_inc_generative_2025. For a recent linguistically-focused comparison of NMT, LLM, and human translation outputs, see sizov-etal-2024-analysing. kong_evaluation_2025 is an example of a comparative user study of the translation performance of ChatGPT-4 and Google Translate in a very practical domain of hospital discharge instructions.
@@ 3.6 Multilingualism
While some recent NMT models are multilingual by design (i.e. trained on parallel sentences in multiple languages and capable of translating between a chosen pair; e.g. NLLB), most are bilingual (one language pair). LLMs, in contrast, tend to be trained on multiple languages as a necessity: even if 90–95% of the data is English, the remaining 5–10% might cover dozens of languages (blevins_language_2022; meta_llama_2025). As a result, LLMs often have broader multilingual scope than a given NMT model (cui_multilingual_2025). A single LLM like GPT-5 or Llama can translate between many languages (English to French, French to German, Japanese to Russian, etc.), whereas an NMT model is usually limited to the directions it was trained on (unless it’s a massively multilingual model like NLLB). This breadth is advantageous, but it comes with uncertainty: the LLM may not be equally strong in all languages. Some languages with very little representation in pretraining data see much lower translation quality. [^5] E.g., zhu_multilingual_2024 note that GPT-4 still struggled on several African and Southeast Asian languages. NMT models can also struggle with low-resource languages, of course, but one can explicitly gather parallel data for those languages to train or fine-tune the model, whereas an LLM’s fixed pretraining mix might under-represent them (unless augmented with further training).
@@ 3.7 Behavior and Errors
NMT and LLM systems also tend to exhibit different failure modes. NMT models, when they err, often produce mistranslations that are locally implausible (e.g. garbled or nonsensical phrases) or they might drop segments of the input (omissions) if attention-based alignment fails. LLMs, by contrast, rarely produce outright garble; thanks to their strong language modeling priors their errors are usually fluent but can be misleading. A common LLM translation error is hallucination—inserting content absent from the source, or over-smoothing—translating idiomatic expressions literally, thus losing nuance. [^6] But he_exploring_2024 report that prompting an LLM to mimic a human translator’s step-by-step process can reduce such deficiencies. LLMs might also be inconsistent in style or level of formality unless instructed, whereas an NMT system can be constrained by training data style or use explicit formality tags. These differences remind us that an LLM is trying to produce a plausible continuation given the prompt and its training, whereas an NMT is laser-focused on reproducing the source content in target form.
@@ 4 Recent Uses of LLMs in Translation Tasks
The application of LLMs to translation-related tasks began almost as soon as LLMs themselves burst onto the scene (brown_language_2020; chowdhery_palm_2022; openai_gpt-4_2023). Researchers and practitioners have tested cutting-edge LLMs on various multilingual tasks, often with impressive results. In this section, I review recent uses of LLMs in translation and other cross-lingual contexts. Key themes include zero-shot vs. few-shot prompting, performance on low-resource languages, pivot translation strategies, chain-of-thought prompting for translation, and exploiting the inherently multilingual nature of latent representations in LLMs.
@@ 4.1 Zero- and Few-Shot Translation with LLMs
One of the headline capabilities of LLMs is zero-shot learning—performing a task without direct task-specific training (brown_language_2020). Translation has become a primary example. Prompting an LLM with a simple instruction “Translate the following sentence to Mandarin:________” can yield a credible translation even if the model was never explicitly trained on parallel data for that language. Early tests with GPT-3 showed surprisingly competent zero-shot translation for several languages (castilho_online_2023), sparking immediate interest in using such models as universal translators (savenkov_gpt-3_2023).
Subsequent research has systematically evaluated zero-shot vs. few-shot prompting for translation (e.g. zhang_machine_2023). Few-shot prompting means providing a few example translations in the prompt before asking the model to translate a new sentence. For instance, one might input:
@@
and let the model continue. This technique often boosts translation accuracy, especially for models not specifically instruction-tuned. garcia_unreasonable_2023 dubbed it the “unreasonable effectiveness of few-shot learning,” finding that even 1–2 examples can significantly improve the translation output output for certain LLMs. The improvement is typically more pronounced for smaller or less instruction-savvy models. On the other hand, for GPT-4 and later models, the gap between zero-shot and few-shot regimes is smaller, since the former is already strong. Later work noted that few-shot translation prompting can, in effect, make up for the lack of translation-specific instruction tuning, or of any instruction tuning.
In fact, curious phenomena have been observed. For example, zhu_multilingual_2024 found that, when given in-context exemplars, LLMs sometimes ignore “unreasonable instructions” asking the model to translate into a different language or do something else altogether, e.g. summarize a sentence instead of translating it. The authors also found that cross-lingual exemplars (in different language pairs than the target pair) can sometimes provide better guidance for low-resource translation than same-pair exemplars. Thus, inserting several RU-EN sentence pairs in the prompt, following by un unrelated Chinese sentence produces a better English translation of it than ZH-EN examples. A hypothetical explanation may be that showing the model examples of, say, RU-EN translation may activate its “translation mode” generally, even if we actually want it to translate from Chinese to English, a direction for which it saw no direct examples. This aligns with the idea that the LLM has a general concept of “translating” that is somewhat language-agnostic. The general lesson appears to be that instruction semantics can be overridden by example context.
@@ 4.2 Low-Resource Languages
For extremely low-resource languages, with virtually no representation in the LLM’s pre-training data, neither pivoting nor prompt optimization may suffice. In such cases, researchers have explored fine-tuning LLMs on supplemental parallel data in a low-resource language pair. The hope is the LLM’s general knowledge transfers, and only minimal data is needed to teach it a new language mapping. Early results show some promise, but the cost of fine-tuning an LLM can be prohibitive. Fortunately, techniques such as LoRA (Low-Rank Adaptation, hu_lora_2021) can be used to inject translation capabilities for new languages into an existing LLM at relatively low compute cost. Thus zhang_machine_2023 used an even more efficient method (QLoRA, dettmers_qlora_2023) to effectively turn a general LLM into a specialized translator with just a fraction of parameters updated, something that can be done on a single consumer-grade GPU.
@@ 4.3 Pivoting and Intermediate Languages
Pivoting (translating through an intermediate language) has been a longstanding strategy in MT for dealing with language pairs lacking direct parallel data. With LLMs, explicit pivoting can be attempted by prompting the model to go through two steps; for example: “Translate this sentence from German to English first; then translate the English output to Hindi.” This prompt can sometimes yield a better Hindi translation than prompting the model to translate directly from German to Hindi (cf. jiao_is_2023).
@@ 4.4 Chain-of-Thought Prompting for Translation
Chain-of-thought (CoT) prompting is a technique where the model is asked to produce intermediate reasoning steps before the final answer (wei_chain--thought_2022). Originally developed for math and logic problems, it has been applied to help LLMs improve their translation output by following, in some cases, a human translation workflow model. For example, briakou_translating_2024 engaged Gemini 1.5 Pro in a multi-turn interaction involving pre-translation research, drafting, refining, and proofreading, which resulted in notable quality improvements.
CoT prompting could be particularly useful when dealing with longer documents. he_exploring_2024 adopted this approach and asked LLMs to generate keywords, topics, and potential translations of tricky terms before attempting the full translation and then used those to guide the process. Since state-of-the-art LLMs have very large context windows (1M+ tokens in some cases), one can feed an entire document and prompt the model to work section by section. Researchers are actively exploring this: for instance, how to prompt an LLM to decide when to translate literally and when to adapt, essentially giving the model a strategy akin to a human translator’s decision-making (jiao_parrot_2023; lu2024mqmapehighqualityerrorannotation). This ventures into the territory of interactive translation with LLMs: rather than using a single prompt, one can craft a sequence of prompts for a genuinely collaborative scenario. There have been experiments using LLMs as translator assistants, where the human can query the model for alternate phrasings, explanations of source text nuances, error annotation, and more (jiao_parrot_2023; ki_guiding_2024; treviso_xtower_2024).
Another interesting recent study develops “compositional translation,” a new method that uses LLMs to improve translation quality for low-resource languages (zebaze_compositional_2025). The approach works by breaking down complex sentences into simpler, semantically coherent phrases, translating each phrase using contextually similar examples, and then merging these translations into a coherent final result. Experiments show that this method can outperform traditional few-shot translation techniques in LLMs, making it a promising solution for low-resource challenges.
@@ 4.5 Multilingual Latent Representations and Cross-Lingual Tasks
Beyond direct translation, LLMs have shown capability in a variety of cross-lingual tasks, such as cross-lingual question answering (answering in English from a French context, etc.), multilingual summarization (summarizing a document in another language), code-switched dialogue understanding, and so on (chua_crosslingual_2023; gao_understanding_2025; blum2025rosettastoneunificationforces). Presumably, all these tasks leverage the model’s multilingual latent space. But what, exactly, is the shape of this space (ravisankar_can_2025)? Where in the model should we look for it? And what methods could be used to study it?